Rita invested 11,000 in two savings accounts. One account pays at a rate of 9% per year and the other pays 11% per year. After one year , she has earned a total of 1,150 in intrest. How much did rita invest in each account?
Answer by stanbon(48554)

Rita invested 11,000 in two savings accounts. One account pays at a rate of 9% per year and the other pays 11% per year. After one year , she has earned a total of 1,150 in intrest. How much did rita invest in each account?
----------
Let amount invested at 9% be "x".
Interest on that is 0.09x dollars
-------
Amount invested at 11% is "11000-x"
Interest on that is 0.11(11000-x)=1210-0.11x dollars
------------
EQUATION:
interest + interest = 1150
0.09x+1210-0.11x = 1150
-0.02x=-60
x=$3000 (amount invested at 9%)
11000-x =$8000 (amount invested at 11%)
Cheers,
Stan H.